OpenStudy (anonymous):

Add (x^2+11 over x^2-5x)+(x^2-8x over x^2-5x)

OpenStudy (anonymous):

\[\frac{2 x^4-10 x^3-8 x+11}{x^2} \]

OpenStudy (anonymous):

(2x^2-8x+11)/x(x-5) if anyone can confirm.. I'm way too tired to do this again ;D

OpenStudy (anonymous):

\[\frac{x^2+11}{x^2-5x}+\frac{x^2-8x}{x^2-5x}\] \[=\frac{2x^2-8x+11}{x^2-5x}\]

OpenStudy (anonymous):

Ok cool.. got the right answer just didn't distribute x to make x^2-5x

OpenStudy (anonymous):

\[\frac{2 x^4-10 x^3-8 x+11}{x^2} \]that I posted above is incorrect. Sorry. @hakkunamatata for one has it right.